OpenStudy (megannicole51):

what method of integration would I use to solve ((x^2)/(1-x^2))dx?

OpenStudy (zzr0ck3r):

\[(1-x^2)=(x-1)(x+1)\]now use partial fraction decomposition. this is the first thing I notice but might not be the best rout...give it a shot

OpenStudy (megannicole51):

i was thinking partial fractions? but idk how you would split them up

OpenStudy (psymon):

Can do u-sub. u = 1-x^2, meaning x^2 = 1-u, which would give you (1-u)/u. Then just split that into two fractions and integrate.

OpenStudy (megannicole51):

hahah wow that was a blonde moment...thank you:D

OpenStudy (zzr0ck3r):

\[(1-x^2)=(1-x)(1+x)\]

OpenStudy (zzr0ck3r):

had to fix my mistake:)

OpenStudy (psymon):

Partial decomp just seems a lot longer to do when you can u-sub xD

zepdrix (zepdrix):

We have to borrow one of the x's for the differential du. So let's be careful doing it this way! :o \[\Large \int\limits \frac{x}{1-x^2}(x\;dx)\]\[\Large u=1-x^2 \qquad\to\qquad -\frac{1}{2}du=(x\;dx)\]\[\Large \int\limits \frac{\sqrt{1-u}}{u}\left(-\frac{1}{2}du\right)\]

OpenStudy (psymon):

Its x^2 on top, zep xD

OpenStudy (zzr0ck3r):

he moved that x

zepdrix (zepdrix):

\[\Large \int\limits \frac{x^2}{1-x^2}dx \quad=\quad \int\limits \frac{x}{1-x^2}(x\;dx)\]

OpenStudy (zzr0ck3r):

x*xdx = x^2dx

zepdrix (zepdrix):

We need one of the x's for the dx. See what happens? :(

OpenStudy (psymon):

didn't even notice obviously. Im dumb, lol.

zepdrix (zepdrix):

Hmm yah, my vote is for Partial Fraction Decomp :D heh

OpenStudy (psymon):

Guess that doesn't make it easier :/

zepdrix (zepdrix):

Ya not so much I guess :D I mean if you remember this integral:\[\Large \int\limits \frac{1}{1-x^2}\;dx \quad=\quad \tanh^{-1} x\] Then fine.. :P but most of us don't, lol.

zepdrix (zepdrix):

Did you ever figure this problem out @megannicole51 ? :o

OpenStudy (megannicole51):

yeah i know how to do those now thank you:)